OpenStudy (mandii):

Solve –2x2 +3x – 9 = 0

OpenStudy (anonymous):

use quadratic formula \[x_{1,2} = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\] first you can multiply both sides by -1, and you get \[2x^2 - 3x + 9 = 0\] so a = 2 b = -3 c = 9 \[x_{1,2} = \frac{-(-3) \pm \sqrt{(-3)^2 - 4(2)(9)}}{2(2)}\] \[x_{1,2} = \frac{3 \pm \sqrt{9 - 72}}{4}\] \[x_{1,2} = \frac{3 \pm \sqrt{-63}}{4}\] \[x_{1,2} = \frac{3 \pm i\sqrt{9}\sqrt{7}}{4}\] \[x_{1,2} = \frac{3 \pm 3i\sqrt{7}}{4}\] therefore \[x_{1} = \frac{3}{4} + \frac{3\sqrt{7}}{4}i\] \[x_{2} = \frac{3}{4} - \frac{3\sqrt{7}}{4}i\]